For the reaction 2 Cu2O (s) + O(g)  →

→ 4 CuO (s), calculate ΔGo (in kJ) at 500 K.


1
Expert's answer
2021-07-14T05:06:13-0400

The moles of the Cu2O


Moles = mass / molar mass


Moles of Cu2O = 91.36 g / 143.091 g per mol = 0.63847 mol Cu2O


Now lets calculate the moles of the O2


1 mol O2 = 2 mol Cu2O


0.63847 mol Cu2O* 1 mol O2 / 2 mol Cu2O = 0.31923 mol O2


Now lets calculate the volume of the O2 using the ideal gas equation


PV= nRT


Where n = moles , P=pressure , T= Kelvin temperature , V= volume in L and R= gas constant (0.08206 L atm per K. mol )


V= nRT/P


V= 0.31923 mol * 0.08206 L atm per K. mol * 298 K / 0.9869 atm


V= 7.91 L atm


Now lets convert L atm to Joules


7.91 LK atm * 101.325 J / 1 L atm = 801 J


Convert Joules to kJ


801 J * 1 kJ / 1000 J = 0.801 kJ


Therefore the amount of work done = 0.801 kJ


Now lets calculate the change in the energy of the system


Delta U= q+W


q= -93.22 kJ


delta U = -93.22 kJ + 0.801 kJ


Delta U = -92.42 kJ

Therefore change in the energy = - 92.42 kJ